# Graph Conditional Flow Matching for Relational Data Generation

### v1

@@ 3.3.3 Implementation Details
In our experiments, we tuned hyperparameters to some extent depending on the dataset. These primarily include neural network parameters, such as the number of hidden units in feedforward layers. The size of the node embeddings is also a key hyperparameter. Tuning this value was important to balance expressiveness and overfitting, as overly large embeddings can lead the model to memorize structures. Across all experiments, we constrained the embedding size to values between 2 and 10. In Appendix B we discuss how the validation loss changes as a function of this hyperparameter for one of the datasets.
@@ 4.1 Experimental Settings
We evaluate our method using SyntheRela (Synthetic Relational Data Generation Benchmark)[21], a recently developed benchmark library for relational database generation. This tool enables comparison of synthetic data fidelity (i.e., similarity to original data) across multiple open-source generation methods and various datasets.
@@
We experiment with six real-world relational datasets: AirBnB [24], Walmart [23], Rossmann [25], Biodegradability [3], CORA [32] and the IMDb MovieLens dataset [17, 18], a commonly used dataset to study graph properties, containing users’ ratings to different movies. The last three dataset have tables with multiple parents. Some of these datasets were subsampled to enable comparison with other methods. More details are provided in Appendix C.
@@
Our primary objective is generating high-fidelity synthetic relational data. To evaluate fidelity, we adopt a discriminator-based approach where an XGBoost classifier [8] (often the preferred classifier for tabular data) is trained to distinguish real from synthetic records. Lower discriminator accuracy indicates higher synthetic data quality, with an accuracy of $0.5$ implying indistinguishability. While this is straightforward for single table datasets, where the input of the discriminator is a single row, this is not for relational data. For relational data, we use the SyntheRela library’s Discriminative Detection with Aggregation (DDA) metric [21], which extends single-table discrimination by enriching the content of the rows of parent tables with aggregate information of its "child" rows. In particular, they add to each row of a parent table the count of children, the mean of real-valued fields of children and the count of unique values of categorical value fields. We believe that discriminator-based metrics are a simple, concise and powerful way to evaluate the fidelity of synthetic data. We compare our method against those present in SyntheRela, that are the leading open-source approaches for relational data generation.
@@ 4.2 Results
Our method generally outperforms all baselines, often by a large margin. Moreover, it is applicable to all relational datasets considered, as it can handle complex schema structures, including tables with multiple parent tables, multiple foreign keys referencing the same table, and missing data. Missing results for some baselines are due to their limitations: ClavaDDPM cannot synthesize CORA and Biodegradability, as it only supports a single foreign-key relation between two tables, REaLTabF does not support tables with multiple parents and SDV fails to synthesize the IMDB dataset due to scalability issues [21]. The performance metrics for other methods are taken from Hudovernik [20], where the SyntheRela library was also used for fidelity evaluation. Variability in the reported results is due to different initialization seeds used both for training and generation.
To assess the impact of the embeddings produced by the GNN, we evaluate the performance of our method when embeddings are ablated. This corresponds to training separate single-table models for each table. The results were negatively affected, though not always dramatically, suggesting that the underlying foreign-key graph, which we re-sample, also plays an important role in achieving high fidelity. Additionally, these results reflect potential limitations of the discriminator used during evaluation.
Table 1: Average accuracy with standard deviation of an XGBoost multi-table discriminator using rows with aggregated statistics. For datasets with multiple parent tables, the highest accuracy was selected. For the CORA dataset, we notice that the simple post-processing step consisting of removing duplicated records from a child table ($\approx 3\%$ of records), allowed us to obtain a performance of $\approx 0.50$.
@@
Hudovernik [20] $0.67\pm 0.003$ $0.83\pm 0.01$ $\mathbf{0.60\pm 0.01}$ $0.64\pm 0.01$ $0.77\pm 0.01$ $0.79\pm 0.04$
ClavaDDPM [36] $\approx 1$ $0.83\pm 0.004$ $0.86\pm 0.01$ $0.74\pm 0.05$
RCTGAN [15] $0.98\pm 0.001$ $0.88\pm 0.01$ $0.73\pm 0.01$ $0.95\pm 0.002$ $0.88\pm 0.01$ $0.96\pm 0.02$
REaLTabF. [43] $\approx 1$ $0.92\pm 0.01$ $\approx 1$
SDV [38] $\approx 1$ $0.98\pm 0.01$ $\approx 1$ $0.98\pm 0.003$ $0.90\pm 0.03$
@@
We evaluated potential privacy leaks in each table, where parent tables were enriched with aggregated information as previously described. For each table, we computed the distance-to-closest-record (DCR) [31, 37, 44] of each synthetic and real record comparing to a hold-out set of real records. We consider a synthetic table to exhibit privacy leakage if the percentage of its DCRs falling below the $\alpha$-percentile of the DCRs of real data is significantly greater than $\alpha$ [30, 5, 39]. Intuitively, this indicates that synthetic records are close to real records more often than expected, suggesting potential privacy risk. As shown in Table 2, when considering the $2\%$ percentile this percentage ($p_{\leq 2\%}$) remains close to the expected value of $2\%$. The privacy score [35], a derived statistic that takes a value of zero (or slightly lower) when no privacy risk is detected and one when all synthetic records are deemed risky, is consistently close to zero, indicating negligible privacy risk.
Table 2: Privacy results for tables having at least 100 records, at least 2 columns (after aggregation) and no more than 10% of real DCRs equal to zero.
@@
AirBnB users 10,000 22 $1.95\%\pm 0.08\%$ $-0.0005\pm 0.001$
Biodegradability molecule 328 6 $0.00\%\pm 0.00\%$ $-0.02\pm 0.000$
@@
Walmart depts 15,047 5 $1.01\%\pm 0.04\%$ $-0.01\pm 0.000$
features 225 12 $1.33\%\pm 1.93\%$ $-0.007\pm 0.020$
@@ 5 Related Works
Early approaches for tabular data generation include Bayesian networks [51], autoregressive models [34] and factor graphs [33], that usually required data to be discretized. Early deep latent-variable models as VAE [26], GAN [13] were later extended to model heterogeneous tabular data [49, 48]. More recently there have been works leveraging transformer-based deep autoregressive model [7] and diffusion models [27], or latent diffusion models [50]. Notably, in Jolicoeur-Martineau et al. [22] they use flow-matching where the denoiser is based on trees, which have often been shown to be state of the art for several tasks relating to tabular data [14, 42, 4].
@@
Synthesizing relational data introduces the additional challenge of preserving inter-table dependencies. The Synthetic Data Vault (SDV) [38] pioneered multi-table synthesis via the Hierarchical Modeling Algorithm (HMA), which employs Gaussian copulas and recursive conditional parameter aggregation to propagate child-table statistics into parents. GAN-based relational extensions include Gueye et al. [15], conditioning child-table synthesis on parent and grand-parent row embeddings, and Li and Tay [28], where the generation of single rows is based on GANs, but tables are generated sequentially in an autoregressive way, following the foreign-key topology. Based on a similar principle Solatorio and Dupriez [43] and Gulati and Roysdon [16] leverage instead a transformer based autoregressive model. These methods still cannot properly manage the generation of tables with multiple parents, since the number of children per parent’s row depends only on one of the parents. In general these sequential methods can only properly deal with tree-like topologies. Pang et al. [36] introduce a guided diffusion approach using latent variables to capture long-range dependencies across tables. To handle tables with multiple parents, they generate a version of the child table for each parent and heuristically merge them by selecting similar rows. In Xu et al. [46] the graph is first generated with a statistical method, and then the content of each table is sequentially generated by conditioning to the content of already generated tables and topological information. A similar approach is proposed in concurrent and independent work by Hudovernik [20], sharing similarities to ours. Data generation is based on latent-diffusion, conditioned on a pre-generated graph by node embeddings encoding topological and neighborhood information, computed using a GNN. Our work differs in three aspects: (1) we employ flow-matching rather than latent diffusion; (2) our GNN is integrated in the denoiser, so it is trained end-to-end, whereas theirs uses embeddings precomputed independently on the generative models of the records; (3) we generate tables in parallel rather than sequentially.
@@ 6 Limitations
Foreign-key Graph Generation.
As this method allows to generate the content of a relational dataset, but not the foreign-key graph, it has to be combined with a graph generation algorithm in order to properly generate novel relational data. However, we think separating the two problems is a promising approach. For example, in Xu et al. [46] they use a statistical method to generate the graph. We follow the same simple approach of Hudovernik [20] for sampling the graphs, that is resampling the original connected components. This could potentially raise a privacy issue, that is however, outside the scope of this work, as we aim at building an effective generative model conditioning on a given foreign-key graph. Nevertheless, we did not observe any privacy leaks in the analysis of parent tables enriched with aggregated information.
Scaling.
Our method requires a GNN to process whole connected components. This is potentially problematic when these are very large. There are many approaches to deal with this as advanced batching strategies, graph partitioning, or out-of-core processing. Nevertheless, in our case it was never an issue, as we were able to fit the whole datasets into memory. So the computational complexity of our method scales linearly with the size of the largest connected component.
@@ 7 Conclusion
We proposed a novel approach for generating relational data, given the graph describing the foreign-key relationships of the datasets. Our method uses flow matching to build a generative model of the whole content of a relational dataset, exploiting a GNN to increase the expressiveness of the denoiser, by letting information flow across connected records. Our method achieves state-of-the-art performance in terms of fidelity of the generated data for several datasets, when compared to performances of other open-source methods present in the SyntheRela benchmark library. Moreover, we did not observe any privacy leakage in the generated synthetic tables, even when parent records were enriched with aggregated statistics from their child tables.
@@
An interesting direction of development is the combination with generative models for large graphs, such as exponential random graphs models [40]. We think this approach is promising as the computational complexity of our method scales linearly with the size of the dataset, while deep generative models of graphs often scale quadratically [52]. Moreover, foreign-key graphs are often simple enough to be modeled by less powerful but scalable statistical models. As our method is based on flow-matching, one can further exploit properties of diffusion-like models, as guidance, inpainting, or generating variations of a given dataset. Finally, we believe there is a margin of improvement of our models within our framework, as we used relatively simple neural network architectures.
Broader Impact.
Our work contributes to the development of high fidelity relational data generation techniques. We believe these can have a positive impact, as synthetic data can enable privacy-preserving sharing of socially valuable information (e.g., medical data). However, malicious use is also possible, for instance, for data counterfeiting.
@@ Appendix B Technical Details on Training, Architecture, and Hyperparameters
We provide here a small overview of the technical details regarding neural network training and architecture, for more details, we refer the reader to the code released at https://github.com/DavideScassola/graph-conditional-flow-matching.
@@ B.2 GNN Architectures
Since the graphs are heterogeneous, with multiple node types and edge types, standard GNN layers must be extended to handle this structure. Following the design pattern in PyTorch Geometric for heterogeneous message passing [12], we define the modified layer as:
@@
Our primary GNN architecture is based on the GATv2 convolution layer [6]. Each GNN block contains a GATv2 convolution followed by a residual linear transformation and ReLU activation:
@@
We also experiment with a variant based on the Graph Isomorphism Network (GIN) [47]. Each node is first projected into a shared latent space using a type-specific linear layer:
@@ B.3 Table-specific Denoisers
Each denoiser $f^{k}$ is implemented as a multi-layer perceptron (MLP), where the input is the concatenation of three components: the current noisy value $\boldsymbol{x}^{i}_{t}$, the time $t$ (embedded following Dhariwal and Nichol [9]), and the node embedding $\varepsilon^{i}$.
We apply Layer Normalization [1] after each hidden activation (SiLU [11]), except in the final layer. The output layer is linear and restores the original dimensionality of $\boldsymbol{x}^{i}_{t}$. For one-hot-encoded categorical features, a final softmax activation is applied.
@@ Appendix C Datasets
We summarize the datasets used in our experiments in Table 4. To allow comparisons with prior work, the AirBnB, Rossmann, and Walmart datasets were downsampled. AirBnB and Rossmann are relatively simple, each containing a single foreign key relationship, while the Walmart dataset features a table with two child tables. The IMDB, Biodegradability, and CORA datasets have more complex structures, including multiple foreign keys and tables referencing multiple parent tables. In particular, both Biodegradability and CORA include tables with two distinct foreign keys pointing to the same parent table. For example, the cites table in CORA contains both the ID of the citing paper and that of the cited paper. Additional details about the datasets are provided in Hudovernik et al. [21].

### v2

@@ 3.3 Generative Modeling from Single-Sample Data
In our experiments, we tuned hyperparameters to some extent depending on the dataset. These primarily include neural network parameters, such as the number of hidden units in feedforward layers. The size of the node embeddings is also a key hyperparameter. Tuning this value was important to balance expressiveness and overfitting, as overly large embeddings can lead the model to memorize structures. Across all experiments, we constrained the embedding size to values between 2 and 10. In Appendix B we discuss how the validation loss changes as a function of this hyperparameter for two of the datasets.
@@ 4.1 Experimental Settings
We evaluate our method using SyntheRela (Synthetic Relational Data Generation Benchmark)(hudovernik2024benchmarking), a recently developed benchmark library for relational database generation. This tool enables comparison of synthetic data fidelity (i.e., similarity to original data) across multiple open-source generation methods and various datasets.
@@
We experiment with six real-world relational datasets: AirBnB (airbnb), Walmart (walmart), Rossmann (rossmann), Biodegradability (biodegradability), CORA (cora) and the IMDb MovieLens dataset (harper2015movielens; movielens), a commonly used dataset to study graph properties, containing users’ ratings to different movies. The last three dataset have tables with multiple parents. Some of these datasets were subsampled to enable comparison with other methods. More details are provided in Appendix C.
@@
Our primary objective is generating high-fidelity synthetic relational data. To evaluate fidelity, we adopt a discriminator-based approach where an XGBoost classifier (chen2016xgboost) (often the preferred classifier for tabular data) is trained to distinguish real from synthetic records. Lower discriminator accuracy indicates higher synthetic data quality, with an accuracy of $0.5$ implying indistinguishability. While this is straightforward for single table datasets, where the input of the discriminator is a single row, this is not for relational data. For relational data, we use the SyntheRela library’s Discriminative Detection with Aggregation (DDA) metric, which extends single-table discrimination by enriching the content of the rows of parent tables with aggregate information of its ”child” rows. In particular, they add to each row of a parent table the count of children, the mean of real-valued fields of children and the count of unique values of categorical value fields. We believe that discriminator-based metrics are a simple, concise and powerful way to evaluate the fidelity of synthetic data. We compare our method against those present in SyntheRela, that are the leading open-source approaches for relational data generation.
@@ 4.2 Results
Our method generally outperforms all baselines, often by a large margin. Moreover, it is applicable to all relational datasets considered, as it can handle complex schema structures, including tables with multiple parent tables, multiple foreign keys referencing the same table, and missing data. Missing results for some baselines are due to their limitations: ClavaDDPM cannot synthesize CORA and Biodegradability, as it only supports a single foreign-key relation between two tables, REaLTabF does not support tables with multiple parents and SDV fails to synthesize the IMDB dataset due to scalability issues (hudovernik2024benchmarking). The performance metrics for other methods are taken from hudovernik2024relational, where the SyntheRela library was also used for fidelity evaluation. Variability in the reported results is due to different initialization seeds used both for training and generation.
To assess the impact of the embeddings produced by the GNN, we evaluate the performance of our method when the embeddings are ablated. This corresponds to training separate single-table models for each table. The results were negatively affected, with the only exception being the CORA dataset. Nevertheless, we observed that ablating the GNN always led to a significant increase in validation loss, thus suggesting potential limitations of the discriminative metric in evaluating certain datasets.
@@
hudovernik2024relational $0.67\pm 0.003$ $0.83\pm 0.01$ $\mathbf{0.60\pm 0.01}$ $0.64\pm 0.01$ $0.77\pm 0.01$ $0.79\pm 0.04$
ClavaDDPM $\approx 1$ - - $0.83\pm 0.004$ $0.86\pm 0.01$ $0.74\pm 0.05$
RCTGAN $0.98\pm 0.001$ $0.88\pm 0.01$ $0.73\pm 0.01$ $0.95\pm 0.002$ $0.88\pm 0.01$ $0.96\pm 0.02$
REaLTabF. $\approx 1$ - - - $0.92\pm 0.01$ $\approx 1$
SDV $\approx 1$ $0.98\pm 0.01$ $\approx 1$ - $0.98\pm 0.003$ $0.90\pm 0.03$
Table 1: Average accuracy with standard deviation of an XGBoost multi-table discriminator using rows with aggregated statistics. For datasets with multiple parent tables, the highest accuracy was selected. The CORA dataset is the only one for which using GNN embeddings does not improve the evaluation metric. However, we noticed that the simple post-processing step consisting of removing duplicated records from a child table ($\approx 3\%$ of records), allowed us to obtain a performance of $\approx 0.50$. Moreover, we observed a lower validation loss when the GNN was used. Statistics are computed over three different runs.
@@
We evaluated potential privacy leaks in each table, where parent tables were enriched with aggregated information as previously described. For each table, we computed the distance-to-closest-record (DCR) (dcr2; dcr3) of each synthetic and real record comparing to a hold-out set of real records. We consider a synthetic table to exhibit privacy leakage if the percentage of its DCRs falling below the $\alpha$-percentile of the DCRs of real data is significantly greater than $\alpha$ (quantile1; quantile3). Intuitively, this indicates that synthetic records are close to real records more often than expected, suggesting potential privacy risk. As shown in Table 2, when considering the $2\%$ percentile this percentage ($p_{\leq 2\%}$) remains close to the expected value of $2\%$. The privacy score (palacios2025contrastive), a derived statistic that takes a value of zero (or slightly lower) when no privacy risk is detected and one when all synthetic records are deemed risky, is consistently close to zero, indicating negligible privacy risk in the content of the synthetic relational data.
@@
AirBnB users 10,000 22 $1.95\%\pm 0.08\%$ $0.0005\pm 0.001$
Biodegradability molecule 328 6 $00\%\pm 0.00\% 0.0$$02\pm 0.000$
@@
Walmart depts 15,047 5 $1.01\%\pm 0.04\%$ $0.01\pm 0.000$
features 225 12 $1.33\%\pm 1.93\%$ $0.007\pm 0.020$
Table 2: Privacy results for tables having at least 100 records, at least 2 columns (after aggregation) and no more than 10% of real DCRs equal to zero. Statistics are computed over three different runs.
@@ 5 Related Works
Early approaches for tabular data generation include Bayesian networks (zhang2017privbayes), autoregressive models (nowok2016synthpop) and factor graphs (mckenna2019graphical), that usually required data to be discretized. Early deep latent‐variable models as VAEs and GANs were later extended to model heterogeneous tabular data (ctgan; tgan). More recently there have been works leveraging transformer-based deep autoregressive model (castellon2023dp) and diffusion models (kotelnikov2023tabddpm; zhang2023mixed; shi2025tabdiff). Notably, in jolicoeur2024generating they use flow matching where the denoiser is based on trees, which have often been shown to be state of the art for several tasks relating to tabular data (grinsztajn2022tree; shwartz2022tabular; borisov2022deep).
@@
Synthesizing relational data introduces the additional challenge of preserving inter‑table dependencies. The Synthetic Data Vault (SDV) (patki2016synthetic) pioneered multi‑table synthesis via the Hierarchical Modeling Algorithm (HMA), which employs Gaussian copulas and recursive conditional parameter aggregation to propagate child‐table statistics into parents. GAN‑based relational extensions include gueye2022row, conditioning child‑table synthesis on parent and grandparent row embeddings, and li2023irg, where the generation of single rows is based on GANs, but tables are generated sequentially in an autoregressive way, following the foreign‑key topology. Following a similar principle solatorio2023realtabformergeneratingrealisticrelational and gulati2023tabmt leverage instead a transformer-based autoregressive model. These methods still cannot properly manage the generation of tables with multiple parents, since the number of children per parent’s row depends only on one of the parents. In general, these sequential methods can only properly deal with tree-like topologies. pang2024clavaddpm introduce a guided diffusion approach using latent variables to capture long-range dependencies across tables. To handle tables with multiple parents, they generate a version of the child table for each parent and heuristically merge them by selecting similar rows. In xu2022synthetic the graph is first generated with a statistical method, and then the content of each table is sequentially generated by conditioning to the content of already generated tables and topological information. A similar approach is proposed in concurrent and independent work by hudovernik2024relational, sharing similarities with ours. Data generation is based on latent-diffusion, conditioned on a pre-generated graph by node embeddings encoding topological and neighborhood information, computed using a GNN. Our work differs in three aspects: (1) we employ flow matching rather than latent diffusion; (2) our GNN is integrated into the denoiser, so it is trained end-to-end, whereas theirs uses embeddings precomputed independently from the generative models of the records; (3) we generate tables in parallel rather than sequentially.
@@ 6 Limitations
Like other generative models, our approach requires dataset-specific hyperparameter tuning, particularly the depth and width of the employed neural networks, to adapt to the training data size and avoid overfitting. Although the architecture we used is relatively simple, we were able to achieve state-of-the-art performance. Therefore, we believe there is significant room for improvement through more sophisticated model design. This work focuses on the method for training a flexible and powerful generative model for relational data, rather than on the specific architecture of the denoiser.
@@ 7 Conclusion
We proposed a novel approach for generating relational data, given the graph describing the foreign-key relationships of the datasets. Our method uses flow matching to build a generative model of the whole content of a relational dataset, exploiting a GNN to increase the expressiveness of the denoiser, by letting information flow across connected records. Our method achieves state-of-the-art performance in terms of synthetic data fidelity across several datasets, outperforming other open-source methods in the SyntheRela benchmark library. Moreover, we did not observe any privacy leakage in the generated synthetic tables, even when parent records were enriched with aggregated statistics from their child tables.
@@
An interesting direction of development is the combination with generative models for large graphs, such as exponential random graph models (robins2007introduction). We think this approach is promising as the computational complexity of our method scales linearly with the size of the dataset, while deep generative models of graphs often scale quadratically (zhu2022survey). Moreover, foreign-key graphs are often simple enough to be modeled by less powerful but scalable statistical models. Finally, as our method is based on flow matching, one can further exploit properties of diffusion-like models, such as guidance, inpainting, or the generation of variations of a given dataset.
Acknowledgements
This research was supported by Aindo, which has funded the PhD of the first author and provided the computational resources. Additionally, this study was carried out within the PNRR research activities of the consortium iNEST funded by the European Union Next-GenerationEU (PNRR, Missione 4 Componente 2, Investimento 1.5 – D.D. 1058 23/06/2022, ECS_00000043).
@@ Appendix B Technical Details on Training, Architecture, and Hyperparameters
We provide here a small overview of the technical details regarding neural network training and architecture, for more details, we refer the reader to the released code.
@@ B.2 GNN Architectures
Since the graphs are heterogeneous, with multiple node types and edge types, standard GNN layers must be extended to handle this structure. Following the design pattern in PyTorch Geometric for heterogeneous message passing (pytorch_geometric_hetero), we define the modified layer as:
@@
Our primary GNN architecture is based on the GATv2 convolution layer (brody2021attentive). Each GNN block contains a GATv2 convolution followed by a residual linear transformation and ReLU activation:
@@
We also experiment with a variant based on the Graph Isomorphism Network (GIN) (xu2018powerful). Each node is first projected into a shared latent space using a type-specific linear layer:
@@ B.3 Table-specific Denoisers
Each denoiser $f^{k}$ is implemented as a multi-layer perceptron (MLP), where the input is the concatenation of three components: the current noisy value $\boldsymbol{x}^{i}_{t}$, the time $t$ (embedded following dhariwal2021diffusion), and the node embedding $\varepsilon^{i}_{t}$.
We apply Layer Normalization (ba2016layer) after each hidden activation (SiLU (elfwing2017silu)), except in the final layer. The output layer is linear and restores the original dimensionality of $\boldsymbol{x}^{i}_{t}$. For one-hot-encoded categorical features, a final softmax activation is applied.
@@ Appendix C Datasets
We summarize the datasets used in our experiments in Table 4. To allow comparisons with prior work, the AirBnB, Rossmann, and Walmart datasets were downsampled. AirBnB and Rossmann are relatively simple, each containing a single foreign key relationship, while the Walmart dataset features a table with two child tables. The IMDB, Biodegradability, and CORA datasets have more complex structures, including multiple foreign keys and tables referencing multiple parent tables. In particular, both Biodegradability and CORA include tables with two distinct foreign keys pointing to the same parent table. For example, the cites table in CORA contains both the ID of the citing paper and that of the cited paper. Additional details about the datasets are provided in hudovernik2024benchmarking.